OpenStudy (anonymous):

Limit problem I am having trouble with... lim (1/x - 1/8)/(x - 8) x approaches 8

OpenStudy (anonymous):

you have to deal with the compound fraction

OpenStudy (anonymous):

\[\frac{1}{x}-\frac{1}{8}=\frac{8-x}{8x}\] is a start

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

then divide by \(x-8\) and get \[\frac{-1}{8x}\] then replace \(x\) by \(8\)

OpenStudy (anonymous):

Im alittle unlcear on this division step....

OpenStudy (anonymous):

ok lets write is all

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

\[\frac{\frac{1}{x}-\frac{1}{8}}{x-8}\] \[=\frac{\frac{8-x}{8x}}{x-8}=\frac{8-x}{8x}\times \frac{1}{x-8}\] then since \(\frac{8-x}{x-8}=-1\) you end with \[\frac{-1}{8x}\]